OpenStudy (anonymous):

Simplify. Please show all work. 4+8 1/2/(1/4)3

OpenStudy (anonymous):

7

OpenStudy (anonymous):

How did you get 7?

jhonyy9 (jhonyy9):

83/8

jhonyy9 (jhonyy9):

8 1/2 =17/2

OpenStudy (anonymous):

So is both of them the answer jhonyy9

jhonyy9 (jhonyy9):

???

OpenStudy (anonymous):

8 1/3= 17/2 is that the correct answer or the firt one you posted or both

jhonyy9 (jhonyy9):

8 1/3 =25/3